Calculate the caloricity of the 450g of chicken soup, which contains 18% of protein, 18% of fats and 0.7% of carbohydrates


1
Expert's answer
2021-03-09T06:05:02-0500

Protein = 450 × (18/100) = 81 cal

Fats = 450× (18/100) = 81 cal

carbohydrates = 450×(0.7/100) = 3.15 cal